Equal masses of iron and sulphur are heated together to form FeS. What fraction of the original mass of excess reactant is left unreacted ? ( Fe = 56, S = 32 )
 
(A). 0.22
(B). 0.43
(C). 0.86
(D). 0.57
 
Please give detailed solution !!!!!!

Grade:12

3 Answers

Arun
5 years ago
Dear student
 
Let equal mass = w
Mole of Fe = w/56
Mole of sulphur = w/32
Now it should form mole of FeS = w/88  = .0113 w
Now w/56 = 0.0178 w
Left = 0.0178 - 0.0113 = 0.0065 
Now reacted = 0.0065 * 88/100 = .572
Hence unreacted = 1 - 0.572 = 0.428 = 0.43
Harish
5 years ago
Hello Sir, 
 
In this step, ‘Now it should form mole of FeS = w/88  = .0113 w’
why it is w/88 and not 2w/88
 
Please reply !
sarah azmi
4 years ago
 
Since same mass therefore excess left is sulphur so the racted part is 57 therfore fractional valie of unreacted sulphur is 43/100= 0.43